OpenStudy (r.bain01):

i had a math equation that started as 5/7m - 3 = 2/7m + 6 ...i worked it up to 3/7m = 9 ... what do i do now?

OpenStudy (anonymous):

You would divide each side by 3/7 to get m! When you are dividing by factions you have to do the opposite. So you would have to mulitply 7/3 * 9 and you get 21 m=21

OpenStudy (r.bain01):

ok right now i have 9 / 3/7 .... explain the process from there please ...i dont quite understand hw u came to 21 ( bt i trust ur answer)

OpenStudy (anonymous):

Okay, so when you are dividing by fractions you have to flip them then multiply! So this would give you 7/3 * 9 when you multiple this out you get 63/3 and that simplifies to 21!

OpenStudy (r.bain01):

do u have time to help me with more ?

OpenStudy (anonymous):

Yup!

OpenStudy (r.bain01):

giv me a sec im getting the homework ...(its not much i swear)

OpenStudy (r.bain01):

ok ... i have an equation ... 4/5x -1 = 1/5 + 5 .... i think i can work it out bt id like to give u my anser to see if im correct when im done

OpenStudy (anonymous):

Alrighty! I'll solve it right now to see what I get

OpenStudy (anonymous):

I got 10

OpenStudy (r.bain01):

ok me 2 (i had to refer bck to the one u helped with) ..... next one ... 3/7x + 5 =5/7x - 1

OpenStudy (anonymous):

I got 21 haha!

OpenStudy (r.bain01):

um ok im stuck at 2/7x = -6 ....wanted to know if the same rule applies like with the other two we did

OpenStudy (r.bain01):

ok i did it ....i got 21 as well....i get scared wen i see negatives sometimes ...can u help with this ? ..... x + 1/2 = -1/3

OpenStudy (anonymous):

So sorry I barely saw your last reply! I'm about to solve the last one now

OpenStudy (anonymous):

x= 1/6

OpenStudy (r.bain01):

y wasnt it negative ?

OpenStudy (anonymous):

Because you have to get a common denominator in order to subtract. so you get 3/6- 2/6 and get 1/6

OpenStudy (anonymous):

:)

OpenStudy (r.bain01):

ok thank u ....thts it for tonite for me ... whenever i have a prob im going to msg u tho ,if u dnt mind , so b prepared at least once a week :)

OpenStudy (anonymous):

Alrighty! I'm usually on here often! So I'll be here to help! :]